# NeuroRVQ: Multi-Scale EEG Tokenization for Generative Large Brainwave Models

### v2

@@ 5 Experiments
We evaluate NeuroRVQ along two dimensions (i) the quality of its tokenizer, compared with the tokenizer of LaBraM, the current state-of-the-art codebook-based method for signal reconstruction; and (ii) the performance of the NeuroRVQ foundation model, compared with other pre-trained models, fine-tuned for downstream classification tasks.
5.1 Experimental Setup
Hyperparameters / Training Process: NeuroRVQ was trained using the settings described in Appendix B. The tokenizer was trained for 100 epochs with $S=4$ temporal branches and 4 $\mathcal{R}$ RVQ codebooks each consisting of 8 single codebooks $\mathcal{V}_{i}\in\mathbb{R}^{8192\times 128}$ and $\lambda_{circle}=0.4$. The foundation model was trained for 50 epochs in the same datasets. The experiments were run on NVIDIA DGX with 4 NVIDIA Tesla V100 GPUs.
Datasets: We have used 13 large-scale EEG datasets, consisting of 12 public and 1 self-collected ($\sim$235 hours of our own self-collected motor data) datasets (we refer the reader to Appendix D). All datasets were resampled to the same sampling frequency of 200Hz.
5.2 Tokenizer Evaluation
The first step involves comparing the tokenization capabilities LaBraM (Jiang et al., 2024) with NeuroRVQ. We performed both (i) in-distribution and (ii) out-of-distribution signal reconstruction comparisons at a per-patch level.
5.2.1 In-Distribution Evaluation
To ensure a fair comparison and to properly evaluate the in-distribution reconstruction capabilities, both architectures (LaBraM and NeuroRVQ) were trained on the same datasets, with the same train-validation split, and identical training regime, as described in Section 5.1 (see Appendix E for the learning curves). Figure 4 illustrates the reconstruction of a single EEG signal/patch using the LaBraM and NeuroRVQ tokenizers.
Figure 4: Per-Band Analysis of Reconstructed EEG signal from LaBraM and NeuroRVQ codebook-based tokenizers. Green lines denote the input EEG signal, orange the reconstructed EEG signal for LaBraM (ours) and red the reconstructed EEG signal for NeuroRVQ.
Table 1 reports the overall reconstruction performance in the validation set across frequency bands. Across all bands, NeuroRVQ achieves orders-of-magnitude lower reconstruction error compared to LaBraM, demonstrating that NeuroRVQ can produce faithful and robust representation of EEG signals across both broadband and band-specific components.
Table 1: In-distribution validation mean squared error (MSE) reconstruction performance across frequency bands for NeuroRVQ and LaBraM tokenizers. Both tokenizers were trained for 100 epochs.
Raw Signal Delta Theta Alpha Beta Gamma
@@ 5.2.2 Out-of-Distribution Evaluation
We further evaluated the generalization capabilities of both NeuroRVQ and LaBraM in reconstructing EEG signals in two datasets unseen during training: a motor dataset High Gamma (Schirrmeister et al., 2017) and a working memory dataset (Pavlov et al., 2022). For LaBraM, we used two variations: (i) the version trained with the same datasets as NeuroRVQ, highlighted as ours; and (ii) the authors’ pre-trained model, highlighted as original. For each dataset, input patches were extracted using the class-related parts of each trial in the dataset.
Tables 2 and 3 show that NeuroRVQ consistently outperforms both variations of LaBraM across all frequency bands, demonstrating the robustness and superior generalization of NeuroRVQ, even when evaluated on EEG signals from entirely unseen datasets.
Table 2: Out-of-distribution mean squared error (MSE) reconstruction performance on memory task across frequency bands for NeuroRVQ and LaBraM tokenizers (original pre-trained and ours).
Raw Signal Delta Theta Alpha Beta Gamma
LaBraM (orig.) 1.985 1.093 0.216 0.191 0.211 0.054
LaBraM (ours) 1.880 2.757 0.361 0.352 0.161 0.031
NeuroRVQ 0.084 0.046 0.006 0.009 0.022 0.010
Table 3: Out-of-distribution mean squared error (MSE) reconstruction performance on motor task across frequency bands for NeuroRVQ and LaBraM tokenizers (original pre-trained and ours).
Raw Signal Delta Theta Alpha Beta Gamma
LaBraM (orig.) 2.004 0.863 0.326 0.400 0.350 0.064
LaBraM (ours) 1.893 3.313 0.669 0.721 0.339 0.049
NeuroRVQ 0.090 0.032 0.010 0.017 0.038 0.009
5.3 Downstream Performance
We evaluated the downstream task performance of fine-tuned NeuroRVQ in comparison to other fine-tuned LBMs, namely NeuroGPT (Cui et al., 2024), CBraMod (Wang et al., 2025), LaBraM (Jiang et al., 2024), EEGPT (Wang et al., 2024), BIOT (Yang et al., 2023) [^1] BIOT could not be tested on Sleep since the benchmark electrodes are missing from the pre-trained model. and MIRepNet (Liu et al., 2025) [^2] As a motor-related foundation model, MIRepNet was tested only on the motor task. All models were evaluated in four downstream classification tasks (Lee et al., 2025): (a) Motor, a four-class movement classification on High Gamma (Schirrmeister et al., 2017) dataset; (b) Memory, a binary working memory classification on Pavlov et al. (2022); (c) Sleep, a 6-class sleep stage classification on Physionet’s Sleep-EDF (Kemp et al., 2000); and (d) Eyes, eyes open vs. closed classification on Physionet’s Motor (Schalk et al., 2004).These tasks capture a diverse range of BCI paradigms and the datasets were specifically chosen for their minimal spurious artifacts, reducing the likelihood of spurious performance during training (Lee et al., 2025).
We applied the same full fine-tuning procedure to all models, training for 20 epochs since performance plateaued beyond that point. For the sleep task, however, we employed early stopping, as the models began to overfit before reaching 20 epochs. For evaluation, we used a consistent 10-fold subject-independent cross-validation across all models and reported balanced accuracy for each task. (For downstream performance comparison between codebook-based LBMs pre-trained on the same datasets, see Appendix F).
Table 4: Comparison of classification balanced accuracy between NeuroRVQ and other LBMs, finetuned on the same datasets using identical training settings ( denotes early stopping). Bold and underlined values indicate best performance and next-best performance respectively (per task or overall).
@@ 5.3 Downstream Performance
Table 4 shows that, when full-finetuned, NeuroRVQ achieves the best or next-best performance in each task, and the best overall performance, measured as the mean accuracy across all tasks. NeuroRVQ has 5.9M parameters in the backbone model and approximately 3,000 parameters in the classification heads (see Appendix G). Yet, it achieves higher mean accuracy than NeuroGPT, which has 79.5M parameters in its backbone; and it also outperforms CBraMod, which has similar backbone parameters to NeuroRVQ but much larger classification heads (approximately 50M parameters; see Appendix G for a detailed comparison of classification head sizes). This combination of strong generalization and compact size highlights the effectiveness of our NeuroRVQ tokenizer.
@@ 6 Discussion
The NeuroRVQ tokenizer demonstrates state-of-the-art reconstruction capabilities on both in-distribution and out-of-distribution signals, as shown in Section 5.2. Importantly, these results cannot be attributed solely to the size of the codebook but also reflect the inductive biases embedded in the tokenizer design together with the specifically designed training process, which are particularly well suited to the characteristics and complex nature of EEG signals. The main contribution of this work is the NeuroRVQ codebook, which can be integrated into larger LBM architectures and may enable further performance gains. Further improvements could also come from incorporating training principles inspired by causal reasoning, such as more targeted temporal and spatial masking strategies (Barmpas et al., 2024a). All these directions are beyond the scope of this work and could be promising avenues for future research.
@@ 7 Conclusion
In this work, we introduce NeuroRVQ, an LBM that leverages a Residual Vector Quantization (RVQ) codebook-based tokenizer. Our main contribution is the design and proper training of the NeuroRVQ tokenizer, which captures low- and high-frequency components of EEG signals, achieves state-of-the-art reconstruction performance and facilitates the development of more effective LBMs. We further showed that NeuroRVQ outperforms existing pre-trained LBMs on various downstream tasks. In summary, NeuroRVQ advances the development of efficient LBMs by providing a high-fidelity tokenizer. The principles underlying NeuroRVQ’s tokenizer can be extended to other biosignals (see Appendix H), paving the way for broader applications of foundation models in the biosignal space.
LLM Declaration
Portions of this manuscript were edited for clarity and grammar with the assistance of a large language model. All scientific content, results and conclusions are the authors’ own.
@@ B.1 Multi-Scale Temporal Encoder Architecture
Our temporal encoder comprises four temporal branches. Each branch applies the following sequence of operations twice in succession (Figure 1c): an 1-D convolution, followed by group normalization with $N=4$ groups and $C=8$ channels, GELU activation, and pooling. Table 5 lists the filter counts, kernel sizes, padding, and pooling parameters for the two sequences, where $x\rightarrow y$ denotes the configuration value(s) used in the first and second sequence, respectively.
@@ Appendix C Number of Codebooks per RVQ
Figure 5: Validation MSE reconstruction loss for the NeuroRVQ tokenizer with different numbers of codebooks (layers) per RVQ.
@@ Appendix D Datasets
We train NeuroRVQ and LaBraM tokenizers on the same datasets with the same train–validation split and an identical training regime (including preprocessing and training steps), as described in §5.1. Figure 6 shows the validation learning curves for both models. NeuroRVQ achieves nearly two orders of magnitude lower reconstruction MSE compared to LaBraM.
Figure 6: Validation learning curves showing reconstruction MSE for NeuroRVQ and LaBraM tokenizers, trained under identical regimes.
@@
We compare the downstream task performance of NeuroRVQ with LaBraM when both models are pre-trained and fine-tuned on the same datasets and with an identical training regime. We use 10-fold subject-independent cross-validation to evaluate models on the four classification benchmarks described in §5.3. Table 8 shows that, when all else is equal, NeuroRVQ achieves 1%–13% higher balanced accuracy than LaBraM on all tasks, demonstrating the effectiveness of our tokenizer in training codebook-based models.
@@
Appendix G Classification Heads: Architecture & Size
Earlier, in Table 4, we compared the downstream classification performance of NeuroRVQ against five state-of-the-art models: CBraMod, BIOT, EEGPT, NeuroGPT and LaBraM. Here, we report the number of parameters in the task-specific classification head for each model. CBraMod and NeuroGPT employ a 3-layer MLP as their classification heads, according to their original implementations, whereas the others use a single layer. The head sizes vary with the number of target classes (Motor has 4, Memory 2, Sleep 6 and Eyes 2); and for CBraMod, also with the number of EEG channels. Table 9 show that NeuroRVQ has relatively low number of parameters compared to other models, both in the backbone and in the classification heads.
@@ Appendix G Classification Heads: Architecture & Size
Model Backbone Motor Head Memory Head Sleep Head Eyes Head
CBraMod 4.9M 50,081,804 40,481,201 73,207,406 41,121,201
NeuroGPT 79.5M 270,756 270,657 270,822 270,657
BIOT 3.2M 1,028 257 - 257
LaBraM 5.8M 804 201 1,206 201
EEGPT 25.7M 260 65 390 65
NeuroRVQ 5.9M 3,204 801 4,806 801
@@ Appendix H Tokenization of Biosignal Modalities
For ECG, we used the PTB-XL dataset (Wagner et al., 2022), resampled at 200Hz. The NeuroRVQ tokenizer was trained using the settings detailed in Appendix B. Training was conducted for 100 epochs with $S=4$ temporal branches and 4 $\mathcal{R}$ RVQ codebooks, each consisting of 8 single codebooks $\mathcal{V}{i}\in\mathbb{R}^{8192\times 128}$, with $\lambda_{circle}=0.4$. Figure 7 shows signal reconstruction for an ECG signal, demonstrating excellent fidelity.
For EMG, we used the emg2pose dataset (Salter et al., 2024), resampled at 1000Hz. The NeuroRVQ tokenizer used the same configuration as above, except that each of the 4 $\mathcal{R}$ RVQ codebooks contained 16 single codebooks $\mathcal{V}_{i}\in\mathbb{R}^{8192\times 128}$ and the kernel sizes approprietely adapted as detailed in Table 10. Figure 8 shows signal reconstruction for an EMG signal, showing high accuracy in preserving its original features.
@@
Figure 7: Per-Band Analysis of Reconstructed ECG signal from NeuroRVQ codebook-based tokenizer. Blue lines denote the input EMG signal and orange the reconstructed ECG signal for NeuroRVQ.
Figure 8: Per-Band Analysis of Reconstructed EMG signal from NeuroRVQ codebook-based tokenizer. Blue lines denote the input EMG signal and orange the reconstructed EMG signal for NeuroRVQ.

### v3

@@ 5 Experiments
Table 1: In-distribution validation and out-of-distribution mean squared error (MSE) reconstruction performance across frequency bands for NeuroRVQ and LaBraM tokenizers (original pre-trained and ours).
In-Distribution Out-of-Dist. (Task) Raw Signal Delta Theta Alpha Beta Gamma
@@
LaBraM (orig.) ✓- (Memory) 1.985 1.093 0.216 0.191 0.211 0.054
LaBraM (ours) ✓- (Memory) 1.880 2.757 0.361 0.352 0.161 0.031
NeuroRVQ ✓- (Memory) 0.084 0.046 0.006 0.009 0.022 0.010
LaBraM (orig.) ✓- (Motor) 2.004 0.863 0.326 0.400 0.350 0.064
LaBraM (ours) ✓- (Motor) 1.893 3.313 0.669 0.721 0.339 0.049
NeuroRVQ ✓- (Motor) 0.090 0.032 0.010 0.017 0.038 0.009
We evaluate NeuroRVQ along two dimensions: (i) the quality of its tokenizer, compared with the tokenizers of LaBraM (Jiang et al., 2024) and BrainOmni (Xiao et al., 2025), recent state-of-the-art codebook-based methods for signal reconstruction; and (ii) the performance of the NeuroRVQ foundation model, compared with other pre-trained open-source models, fine-tuned for downstream tasks.
5.1 Experimental Setup
Hyperparameters / Training Process: NeuroRVQ was trained using the settings described in Appendix B. The tokenizer was trained for 100 epochs with $S=4$ temporal branches (selected heuristically based on neuroscientific findings  (Newson and Thiagarajan, 2018) - the smallest number that still enables effective disentanglement of the canonical EEG frequency bands) and 4 $\mathcal{R}$ RVQ codebooks each consisting of 8 single codebooks $\mathcal{V}_{i}\in\mathbb{R}^{8192\times 128}$ (see Appendix C) and $\lambda_{circle}=0.4$. The foundation model was trained for 50 epochs in the same datasets. The experiments were run on NVIDIA DGX with 4 Tesla V100 GPUs.
Datasets: We have used 13 large-scale EEG datasets, consisting of 12 public datasets and 1 self-collected motor dataset (we refer the reader to Appendix D). All datasets were resampled to the same sampling frequency of 200Hz.
5.2 Tokenizer Evaluation
We compare signal reconstruction errors at a per-patch level of NeuroRVQ’s tokenizer with LaBraM (Jiang et al., 2024) (in- and out-of-distribution settings) and BrainOmni (Xiao et al., 2025) (out-of-distribution setting).
5.2.1 In-Distribution Evaluation
To ensure a fair comparison and to properly evaluate the in-distribution reconstruction capabilities, both LaBraM and NeuroRVQ were trained on the same datasets, with the same train-validation split, and identical training regime, as described in Section 5.1 (see Appendix E for the learning curves). Figure 4 illustrates the reconstruction of a single EEG signal/patch using the LaBraM and NeuroRVQ tokenizers. Table 1 reports the overall reconstruction performance in the validation set across frequency bands. Across all bands, NeuroRVQ achieves orders-of-magnitude lower reconstruction error compared to LaBraM, demonstrating that NeuroRVQ can produce faithful and robust representations of EEG signals across both broadband and band-specific components.
Figure 4: Per-band analysis of a sample reconstructed EEG signal (from the validation set) from LaBraM and NeuroRVQ codebook-based tokenizers. Green lines denote the input EEG signal, orange the reconstructed EEG signal for LaBraM (ours) and red the reconstructed EEG signal for NeuroRVQ.
@@ 5.2.2 Out-of-Distribution Evaluation
Figure 5: Two sample reconstructed EEG signals from BrainOmni and NeuroRVQ codebook-based tokenizers. Blue lines denote the input EEG signal, green the reconstructed EEG signal for BrainOmni and red the reconstructed EEG signal for NeuroRVQ.
Table 2: Comparison of classification balanced accuracy between NeuroRVQ and other LBMs, finetuned on the same datasets using identical training settings ( denotes early stopping). Bold and underlined values indicate best performance and next-best performance respectively (per task or overall). Tasks follow the (Lee et al., 2025) benchmark for EEG foundation models.
@@
We further evaluated the generalization capabilities of both NeuroRVQ and LaBraM in reconstructing EEG signals in two datasets unseen during training: a motor dataset High Gamma (Schirrmeister et al., 2017) and a working memory dataset (Pavlov et al., 2022). For LaBraM, we used two variations: (i) the version trained with the same datasets as NeuroRVQ, highlighted as ours; and (ii) the authors’ pre-trained model, highlighted as original. For each dataset, input patches were extracted using the class-related parts of each trial in the dataset. Tables 1 show that NeuroRVQ consistently outperforms both variations of LaBraM across all frequency bands, demonstrating the robustness and superior generalization of NeuroRVQ, even when evaluated on EEG signals from entirely unseen datasets.
While LaBraM’s codebook is inspired by VQ-VAE (Esser et al., 2020), BrainOmni adopts a single RVQ codebook discretization inspired by EnCodec (Défossez et al., 2023). NeuroRVQ is also RVQ-based, but differs fundamentally in that we employ multiple (per frequency scale) RVQ codebooks, rather than a single RVQ. We compare the reconstruction generalization capabilities of NeuroRVQ with BrainOmni. We use BrainOmni’s sample data (band-pass filtered between 0.5 and 45Hz) because their model requires device layout information. Since the two architectures employ different signal normalization strategies, we compare them using both strategies to ensure fairness. Table 3 and Figure 5 show that NeuroRVQ consistently outperforms BrainOmni in both normalization schemes.
Table 3: Out-of-distribution mean squared error (MSE) reconstruction performance for NeuroRVQ and BrainOmni tokenizers with both normalization (Norm) schemes in BrainOmni’s sample data.
NeuroRVQ-Norm. BO-Norm.
BrainOmni (BO) 0.450 0.332
NeuroRVQ 0.191 0.275
5.3 Downstream Performance
(a) EMG Modality
(b) ECG Modality
Figure 6: Reconstructed sample biosignals (from the validation sets) from NeuroRVQ codebook-based tokenizer. Blue lines denote the input biosignal and orange the reconstructed biosignal for NeuroRVQ.
We evaluated the downstream task performance of fine-tuned NeuroRVQ in comparison to other fine-tuned LBMs, namely NeuroGPT (Cui et al., 2024), CBraMod (Wang et al., 2025), LaBraM (Jiang et al., 2024), EEGPT (Wang et al., 2024), BIOT (Yang et al., 2023) [^1] BIOT could not be tested on the Sleep task since the benchmark electrodes are missing from the pre-trained model. and MIRepNet (Liu et al., 2025) [^2] As a motor-related foundation model, MIRepNet was tested only on the motor task. All models were evaluated in five downstream classification tasks (Lee et al., 2025): (a) Motor, a four-class movement classification on High Gamma (Schirrmeister et al., 2017) dataset (14 subjects); (b) ERP (Event-Related Potential), a two-class ERP classification on the ERP paradigm from Korean University (Kim et al., 2019) dataset (54 subjects); (c) Memory, a binary working memory classification on (Pavlov et al., 2022) (65 subjects); (d) Sleep, a 6-class sleep stage classification on Physionet’s Sleep-EDF (Kemp et al., 2000) (78 subjects); and (e) Eyes, eyes open vs. closed classification on Physionet’s Motor (Schalk et al., 2004) (103 subjects). This benchmark (Lee et al., 2025) was selected because its tasks capture a diverse range of BCI paradigms with datasets specifically chosen for their minimal spurious artifacts, reducing the likelihood of spurious performance during training.
We applied the same full fine-tuning procedure to all models, training for 20 epochs since performance plateaued beyond that point. For the sleep task, however, we employed early stopping, as the models began to overfit before reaching 20 epochs. For evaluation, we used a consistent 10-fold subject-independent cross-validation across all models and reported balanced accuracy for each task (for downstream performance comparison between codebook-based LBMs— namely, LaBraM and NeuroRVQ— pre-trained on the same datasets see Appendix F). The variance between folds is also reported to highlight model’s robustness.
Table 2 shows that, when full-finetuned, NeuroRVQ achieves the best or next-best performance in each task, and the best overall performance, measured as the mean balanced accuracy across all tasks. NeuroRVQ has 5.9M parameters in the backbone model and approximately 3K parameters in the classification heads. Yet, it achieves higher mean performance than NeuroGPT, which has 79.5M parameters in its backbone; and it also outperforms CBraMod, which has similar backbone parameters to NeuroRVQ but much larger classification heads, approximately 50M parameters (see Appendix G for a detailed comparison of classification head sizes). This combination of strong generalization and compact size highlights the effectiveness of NeuroRVQ’s tokenizer.
@@ 6 Discussion
The NeuroRVQ tokenizer demonstrates state-of-the-art reconstruction capabilities on both in- and out-of-distribution signals (Section 5.2). Importantly, these results cannot be attributed solely to the number of codebooks (single or multiple), their size or their type (e.g., RVQ-based). Instead, they arise from the inductive biases embedded in the tokenizer design which, together with the newly introduced unit circle phase-aware training loss, are particularly well-suited for capturing the complex nature of EEG signals.
The main contribution of this work is the NeuroRVQ codebook, which can be (i) applied to biosignal modalities beyond EEG and (ii) integrated into larger LBM architectures, enabling further performance improvements. Beyond EEG, the NeuroRVQ tokenizer is readily adaptable to other biosignal modalities by adjusting the kernel sizes to capture modality-specific frequency ranges and by scaling the number of RVQ codebooks to model signals of varying complexity. Initial results on EMG and ECG (Appendix H) demonstrate strong reconstruction performance, as illustrated in Figures 6(a) and  6(b) respectively. Further improvements can come from larger LBM architectures and incorporating causal reasoning-inspired training principles, such as more targeted temporal and spatial masking strategies (Barmpas et al., 2024a). These directions are beyond the scope of this work and promising for future research.
@@ 7 Conclusion
In this work, we introduce NeuroRVQ, an LBM that leverages an RVQ codebook-based tokenizer. Our main contribution is the design and training of the tokenizer, which captures low- and high-frequency components of EEG signals, achieves state-of-the-art reconstruction performance and facilitates the development of more effective LBMs. We further showed that NeuroRVQ outperforms existing pre-trained LBMs in various downstream tasks. In summary, NeuroRVQ advances the development of efficient LBMs by providing a high-fidelity tokenizer, while its principles can be extended to other biosignals, paving the way for broader use of foundation models in the biosignal space.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ B.1 Multi-Scale Temporal Encoder Architecture
Our temporal encoder comprises four temporal branches. Each branch applies the following sequence of operations twice in succession (Figure 1c): a 1-D convolution, followed by group normalization with $N=4$ groups and $C=8$ channels, GELU activation, and pooling. Table 4 lists the filter counts, kernel sizes, padding, and pooling parameters for the two sequences, where $x\rightarrow y$ denotes the configuration value(s) used in the first and second sequence, respectively.
@@ Appendix C Number of Codebooks per RVQ
Figure 7: Validation mean squared error (MSE) reconstruction loss for the NeuroRVQ tokenizer with different numbers of codebooks (layers) per RVQ codebook.
@@ Appendix D Datasets
We train NeuroRVQ and LaBraM tokenizers on the same datasets with the same train–validation split and an identical training regime (including preprocessing and training steps), as described in §5.1. Figure 8 shows the validation learning curves for both models. NeuroRVQ achieves nearly two orders of magnitude lower reconstruction mean squared error (MSE) compared to LaBraM.
Figure 8: Validation learning curves showing reconstruction mean squared error (MSE) for NeuroRVQ and LaBraM tokenizers, trained under identical regimes.
@@
We compare the downstream task performance of NeuroRVQ with LaBraM when both models are pre-trained and fine-tuned on the same datasets and with an identical training regime. We use a 10-fold subject-independent cross-validation to evaluate models on four classification benchmarks (Motor, Memory, Sleep, Eyes) as described in §5.3. Table 7 shows that, when all training settings are equal, NeuroRVQ achieves 1%–13% higher balanced accuracy than LaBraM on all tasks, demonstrating the effectiveness of our tokenizer in training codebook-based models.
@@
Appendix G Classification Head Sizes
Previously, in Table 2, we compared the downstream classification performance of NeuroRVQ against other pre-trained state-of-the-art models: CBraMod, BIOT, EEGPT, NeuroGPT and LaBraM. Here, we report the number of parameters in the task-specific classification head for each model. CBraMod and NeuroGPT employ a 3-layer MLP as their classification heads, according to their original implementations, while the others use a single layer. The head sizes vary with the number of target classes (Motor has 4, ERP has 2, Memory 2, Sleep 6 and Eyes 2); and for CBraMod, also with the number of EEG channels. Table 8 shows that NeuroRVQ has a relatively low number of parameters compared to other models, both in the backbone and in the classification heads.
@@ Appendix G Classification Head Sizes
Model Backbone Motor Head ERP Head Memory Head Sleep Head Eyes Head
NeuroGPT 79.5M 270,756 270,657 270,657 270,822 270,657
CBraMod 4.9M 50.1M 40M 40.5M 73M 41M
BIOT 3.2M 1,028 257 257 - 257
LaBraM 5.8M 804 201 201 1,206 201
EEGPT 25.7M 260 65 65 390 65
NeuroRVQ 5.9M 3,204 801 801 4,806 801
@@ Appendix H Tokenization of Biosignal Modalities
For ECG, we used the PTB-XL dataset (Wagner et al., 2022), resampled at 200Hz. The NeuroRVQ tokenizer was trained using the settings detailed in Appendix B. Training was conducted for 100 epochs with $S=4$ temporal branches and 4 $\mathcal{R}$ RVQ codebooks, each consisting of 8 single codebooks $\mathcal{V}{i}\in\mathbb{R}^{8192\times 128}$, with $200ms$ patch time window and $\lambda_{circle}=0.4$. Figure 6(b) shows the reconstruction of an ECG signal, demonstrating excellent fidelity.
For EMG, we used the emg2pose dataset (Salter et al., 2024), resampled at 1000Hz. The NeuroRVQ tokenizer used the same configuration as above, except that each of the 4 $\mathcal{R}$ RVQ codebooks contained 16 single codebooks $\mathcal{V}_{i}\in\mathbb{R}^{8192\times 128}$ (since the complexity of EMG signals and the frequencies of interest are larger than ECG), with $200ms$ patch time window and $\lambda_{circle}=0.4$ and the kernel sizes appropriately adapted as detailed in Table 9. Figure 6(a) shows the reconstruction of an EMG signal, showing high accuracy in preserving its original features.